OpenStudy (pulsified333):

f(t)=7He^t+6L(ln(t)) where H and L are constants. What is the derivative?

OpenStudy (pulsified333):

@mathmale

OpenStudy (mathmale):

Note carefully that you have 2 capital letters, H and L, which represent constants. Treat them in the same manner as you would treat numerals such as 29, 13, 2, etc. f(t)=7He^t+6L(ln(t)) in simplest form would look like f(t)=e^t - ln t. Find the derivative of this new f(t) now, please, for practice.

OpenStudy (mathmale):

f '(t) = ?

OpenStudy (pulsified333):

f'(t)=e^t+1/t

OpenStudy (mathmale):

Perfect. Now, differentiate 10e^t-4.5ln t

OpenStudy (pulsified333):

where did you get 10 and -4.5

OpenStudy (mathmale):

Strictly examples. Just like those mysterious L and H.

OpenStudy (mathmale):

Intentionally exposing you to new situations so that they'll be familiar in the future.

OpenStudy (pulsified333):

10e^t-(9/2t)

OpenStudy (mathmale):

Your 10e^t is fine. But your (9/2t) is ambiguous. How would you fix that? (d/dt) [-4.5 ln t] = ?? Hint: Write (-4.5) first. Then, differentiate ln t. Lastly, multiply these 2 results together.

OpenStudy (pulsified333):

-4.5/t

OpenStudy (mathmale):

Much better. Clear, accurate. Thx again for your perseverance.

OpenStudy (pulsified333):

:D

OpenStudy (mathmale):

Differentiate: f(t)=7He^t+6L(ln(t))

OpenStudy (pulsified333):

7e^(t)+(6/t)?

OpenStudy (mathmale):

What hap to the 'H?'

OpenStudy (pulsified333):

don't constants come out of the equation?

OpenStudy (mathmale):

Not in cases like this: That H and that L are constant coefficients. The H stays with the first term and the L stays wtih the second: f(t)=7He^t+6L(ln(t)) Try again. Hint: 7H is a constant coeff.

OpenStudy (pulsified333):

(7He^t+(6L/t))

OpenStudy (mathmale):

Examples: (d/dt)[Ht]=H but (d/dt) [H] = 0. Why the difference? Yes, you're right. Be sure you label your result. What is the name of the original function you were supposed to differentiate?

OpenStudy (pulsified333):

f'(t)

OpenStudy (mathmale):

Right. So now you have a completely correct result, with label, with H and L treated like the constant coeffs they are. Very good.

OpenStudy (mathmale):

You have f '(t) = (7He^t+(6L/t)) I would first omit the first and last parentheses: 7He^t (no left parenthesis) AND (6L)(1/t) (Your parentheses in (6L/t) are in the wrong place; this expression should look like (6L)(1/t) or simply (6L)/t or even 6L/t.

OpenStudy (mathmale):

Gotta get off the 'Net. Looking forward to working with you again. ;)